mikewwe13:

Function f is an exponential function. By what factor does the output value increase as each input value increases by 1? Enter your answer in the box. ______________

mikewwe13:

1 attachment
celticcat:

let f(x) be ab^x when x = 0 f(x) = 1/16 therefore:- 1/16 = ab^0 = a so a = 1/16 using x = 2 and f(x) = 1/4 we have 1/4 = 1/16 * b ^2 b^2 = 16/4 = 4 b = 2 then 2 is the required factor as each input value increases by 1.